The perimeter of a rectangle is 32 feet find the dimensions if the length is 4 feet longer than three times the width then find the area of the rectangle
Answer by tiffany222(56)

Let width be x
Let length be y
y = 3x + 4

P = (2 * width) + (2 * length)
P = 2x + 2y
32 = 2x + 2(3x + 4)
32 = 2x + 6x +8
32 = 8x + 8
24 = 8x
x = 3
y = 3x + 4
= 3 (3) + 4
= 9 + 4
= 13

A = width * length
A = x * y
A = 3 *13
A = 39

Final Answer:
Area = 39 ft^2